Televisions are sold by the length of the diagonal across the screen. If a new 48-in television screen is 42 in. wide, how tall is the screen to the nearest inch?
Answer by rfer(16322)

42^2+x^2=48^2
1764+x^2=2304
x^2=2304-1764
x^2=540
x=23.24